# Distilling 3D distinctive local descriptors for 6D pose estimation

### v2

@@ Abstract
Three-dimensional local descriptors are crucial for encoding geometric surface properties, making them essential for various point cloud understanding tasks. Among these descriptors, GeDi has demonstrated strong zero-shot 6D pose estimation capabilities but remains computationally impractical for real-world applications due to its expensive inference process. Can we retain GeDi’s effectiveness while significantly improving its efficiency? In this paper, we explore this question by introducing a knowledge distillation framework that trains an efficient student model to regress local descriptors from a GeDi teacher. Our key contributions include: an efficient large-scale training procedure that ensures robustness to occlusions and partial observations while operating under compute and storage constraints, and a novel loss formulation that handles weak supervision from non-distinctive teacher descriptors. We validate our approach on five BOP Benchmark datasets and demonstrate a significant reduction in inference time while maintaining competitive performance with existing methods, bringing zero-shot 6D pose estimation closer to real-time feasibility. Project Website: https://tev-fbk.github.io/dGeDi/
@@ I Introduction
Three-dimensional descriptors (or features) encode geometric surface properties through compact numerical representations [1]. They are essential for various downstream tasks in point cloud understanding, including registration [2], object 6D pose estimation [3], segmentation [4, 5], and navigation [6]. These descriptors can be broadly categorized into local and global based on the information they encode. Local descriptors operate at the point or patch level, capturing fine-grained geometric details [7, 8, 2, 9]. Global descriptors also provide point-level representations, but aggregate geometric information from the entire point cloud [10, 11, 12]. Global descriptors are generally more computationally efficient [13] but often sacrifice generalization across different data domains [2].
Local descriptors can achieve rotation invariance through local reference frames (LRF) [2] or point-pair features (PPF) [14]. LRF-based methods establish a local coordinate system for each point using surrounding geometric information [15]. PPF-based methods compute antisymmetric 4D features from pairs of 3D points and their surface normals [16]. Among local descriptors, GeDi [2] has been successfully applied to 6D pose estimation1116D pose estimation involves determining the rotation and translation of a query object (e.g., CAD model) relative to the reference frame of a target object in a scene (e.g., captured by an RGB-D camera). through the FreeZe method [3] in a zero-shot setting, achieving performances that, for the first time in the literature, nearly match those of state-of-the-art fully-supervised methods. However, GeDi’s inference time is impractical for real-world applications due to its computationally expensive process: LRF calculation[15] for each point followed by a PointNet++-based descriptor encoding step [17] (top Fig. 1).
@@
In this work, we explore how to address this inefficiency by using knowledge distillation [19] to train a student model (bottom Fig. 1) that regresses GeDi local descriptors, i.e., the teacher model. However, this presents several challenges. Distillation in this context requires large-scale point cloud datasets [20, 21, 22]. The rotation and scale invariance properties of the teacher model have to be effectively transferred to the student model. Low inlier ratios must be handled, as only a small subset of descriptors is typically reliable, resulting in a significant imbalance during distillation, where most points provide unreliable supervision. Despite these challenges, our approach presents unique opportunities. A more efficient descriptor extractor could bring zero-shot 6D pose estimation closer to real-time performance, making it more viable for practical applications. Moreover, by tailoring the distillation process specifically for 6D pose estimation, we can improve descriptor robustness to occlusions and partial observations. To the best of our knowledge, this is the first attempt to apply distillation to 3D local descriptors.
@@
To this end, we introduce a distillation approach to train an efficient neural network (student) based on PointTransformerV3 (PTV3) [18] to generate GeDi descriptors (teacher). We refer to our approach as dGeDi, derived from distilled GeDi. A key challenge when distilling GeDi is balancing efficiency and practicality. On the one hand, a naive online distillation approach, where the teacher is computed jointly with the student, is impractical due to GeDi’s slow run-time, making training excessively slow. On the other hand, a naive offline distillation, where teacher features are precomputed, would be highly storage-inefficient. To overcome these limitations, we propose a novel learning via correspondences paradigm. Offline, we precompute features from a query object, while online, we transfer them to the target object present in the scene using ground-truth transformations from synthetic 6D pose estimation datasets. This allows scalable training on large datasets while keeping compute and storage requirements manageable. Another challenge is that, unlike regression, where supervision comes from precise ground-truth annotations, our distillation supervision can include non-distinctive descriptors (e.g., on the BOP Benchmark [23], only 3% of points can be matched through nearest neighbor search in the teacher feature space). Standard loss functions [24, 25] tend to overfit these outliers, producing unreliable student features. To mitigate this, we introduce a novel loss function that weights each point’s contribution based on its registration error. Points with errors above a certain threshold contribute almost equally, preventing excessive penalization from outliers. Since training data are synthetic, we incorporate data augmentation to enhance rotational robustness in the student model and reduce the synthetic-to-real domain gap. We validate dGeDi through direct comparison with GeDi [2] and further evaluate its effectiveness by integrating it into FreeZe [3], a state-of-the-art zero-shot pose estimator. Additionally, we conduct an extensive ablation study to analyze dGeDi’s key components. Experimental results demonstrate that dGeDi significantly reduces inference time while maintaining competitive performance with alternative methods. In summary, our main contributions are:
@@ II-A Zero-shot object 6D pose estimation
Zero-shot pose estimation methods estimate the 6D pose of objects that are unseen during training. These methods train on synthetic dataset or leverage the knowledge acquired by pre-trained vision and geometric foundational models. The first category of methods like MegaPose [26] and ZeroPose [27] uses render and compare strategy. At first a classification network selects the best-matching template from a set of rendered images that closely resemble the input image. Then, a refiner network iteratively determines the pose by matching the selected template with additional viewpoints. Since these method requires a separate inference for each input image and template pair, the computational cost and inference time scale linearly with the number of templates. Methods like GigaPose [28] addresses the issue of inference time scaling. It renders templates to extract dense features using a Vision Transformer (ViT), and then finds the best template using fast nearest neighbour search in the feature space. On top, two lightweight MLPs estimate the 2D scale and in plane rotation from a single 2D-2D correspondence and the two relative patch features extracted by another feature extractor. SAM-6D [29] leverages ViT to extract features that are back-projected into 3D to find 3D-3D correspondences between two point clouds. SAM-6D uses a matching network to enhance features before finding the actual correspondences. FoundationPose [30] propose a generalizable and unified pipeline for pose tracking. The coarse pose is estimated using a trained transformer based network. This is followed by a pose refinement network that operates on input image crops and the renderings of the object model. The refined poses are then ranked using a self attention based transformer network.
Other methods do not require task-specific training and leverage the knowledge learned by large scale vision and geometric foundation models. ZS6D [31] extracts features using pre-trained DINOv2 from the inputs image and templates. Then, the best template is chosen using the cosine similarity and 2D-3D correspondences are retrieved using patch-wise similarity. The pose is estimated using PnP and RANSAC. FoundPose [32] integrates DINOv2 patch features into bag-of-words representation for efficient template retrieval. The pose hypothesis is generated using PnP and RANSAC and is refined using optimization-based photometric alignment algorithm. FreeZe [3] fuses descriptors from pre-trained geometric and vision foundation models to extract discriminative features. This is followed by RANSAC based registration and a symmetry-aware pose refinement.
@@ III Preliminaries: FreeZe method
Preprocessing. Let $\mathcal{P}^{Q}$ and $\mathcal{P}^{T}$ be the 3D point cloud of $Q$ and $T$, respectively. $\mathcal{P}^{Q}$ is obtained by sampling points on $\mathcal{O}$’s surface. $\mathcal{P}^{T}$ is obtained by using an off-the-shelf zero-shot segmentor [29] to extract the crop $^{T}\subset\textbf{I}$ I containing $T$, and then lifting $\textbf{I}^{T}$ in 3D using sensor’s intrinsic parameters.
Feature extraction. FreeZe extracts features $\mathcal{F}^{Q}$, $\mathcal{F}^{T}$ from $\mathcal{P}^{Q},\mathcal{P}^{T}$ using frozen vision [53] and geometric [2] encoders, i.e., $\mathcal{F}=(\mathcal{F}_{\text{geo}}|\mathcal{F}_{\text{vis}})$, where $\mathcal{F}_{\text{geo}}$ are the geometric features, $\mathcal{F}_{\text{vis}}$ are the visual features, and |||| is the concatenation operator.
@@ IV Object-oriented distillation
While FreeZe performs feature extraction with both geometric and vision encoders, we focus on the distillation of former ones222 Object-oriented distillation can also be applied to the vision encoder, but it is out of scope for this work., assuming $\mathcal{F}=\mathcal{F}_{\text{geo}}$. We introduce a novel distillation procedure to transfer the distinctive properties of a slow teacher encoder to a faster student encoder (Fig. 2). We define our proposed distillation approach as object-oriented because training supervision is provided at the object level. First, we extract teacher descriptors $\mathcal{F}^{Q}=\Phi_{\Theta}(\mathcal{P}^{Q}),\mathcal{F}^{T}=\Phi_{\Theta}(\mathcal{P}^{T})$ using GeDi [2] encoder $\Phi_{\Theta}$. Then, we learn student descriptors $\mathcal{G}^{Q}=\Psi_{\Omega}(\mathcal{P}^{Q}),\mathcal{G}^{T}=\Psi_{\Omega}(\mathcal{P}^{T})$ with a PTV3 [18] encoder $\Psi_{\Omega}$. During distillation, we optimize the parameters $\Omega$ so that $\mathcal{G}^{Q}\approx\mathcal{F}^{Q}$ and $\mathcal{G}^{T}\approx\mathcal{F}^{T}$, while $\Theta$ remain frozen. Note that this differs from online knowledge distillation approaches [56], where both the teacher and student are neural networks and learn $\Theta,\Omega$ simultaneously. In our case, the architectures of $\Phi_{\Theta}$ and $\Psi_{\Omega}$ are distinct (Fig. 1). Teacher features are precomputed, stored in memory, and loaded as needed during distillation. To optimize this process, we propose storing teacher features only for query objects and introduce a module that leverages ground-truth 6D poses to transfer them to target objects (Sec. IV-A). Moreover, we propose a custom loss function that focuses learning on noise-free points, leading to improved performance (Sec. IV-B).
@@ IV-A Learning via correspondences
During training, given a data sample $(\mathcal{P}^{Q},\mathcal{P}^{T},\textbf{T}^{Q\to T})$, we leverage the ground-truth 6D pose $\textbf{T}^{Q\to T}$ to align $\mathcal{P}^{Q}$ with $\mathcal{P}^{T}$ and perform a nearest neighbor search in the metric space to establish the correspondence $\Gamma\colon\mathcal{P}^{T}\to\mathcal{P}^{Q}$, where each point $\textbf{p}\in\mathcal{P}^{T}$ is mapped to $\textbf{q}=\operatorname*{argmin}_{\textbf{q}\in\mathcal{P}^{Q}}\lVert\textbf{T}^{Q\to T}\textbf{q}-\textbf{p}\rVert$, where p and q represent the 3D coordinates of corresponding points. We use $\Gamma$ to transfer query teacher descriptors $\mathcal{F}^{Q}=\Phi_{\Theta}(\mathcal{P}^{Q})$ to $\mathcal{P}^{T}$, obtaining ~$\tilde{\mathcal{F}}^{T}=\Gamma(\mathcal{F}^{Q})$. By replacing the explicit computation of $\mathcal{F}^{T}=\Phi_{\Theta}(\mathcal{P}^{T})$ with ~$\tilde{\mathcal{F}}^{T}$ during distillation, we achieve two key benefits: (i) improved supervision for target objects and (ii) a scalable, object-oriented distillation process.
Improved target supervision. Matching features $\mathcal{F}^{Q},\mathcal{F}^{T}$ is challenging because input data differ significantly: $\mathcal{P}^{Q}$ is computed from noise-free synthetic CAD models, while $\mathcal{P}^{T}$ is extracted from noisy real-world RGBD images. This domain gap results in variations in point densities, partiality, and artifacts created by lifting the input image I to generate $\mathcal{P}^{T}$. For example, on the MegaPose-GSO [26] dataset, we measured that, on average, only 21.4% of the points can be matched via nearest neighbor search in the teacher feature space. This means that 78.6% of the features are considered outliers for feature matching. In contrast, by transferring query teacher features to target point clouds with $\Gamma$, we simulate the ideal scenario where all target points are matched with the corresponding query points, promoting reliable supervision during distillation.
@@ IV-B Loss function design
Although GeDi has proven to be both generalizable and distinctive [2], it still exhibits low inlier ratios, with an average RON of 3% on the BOP Benchmark and 21.4% on MegaPose-GSO. This indicates that only a small fraction of points are highly distinctive, while the majority have ambiguous features. Consequently, a loss function that assigns equal weight to all points will be dominated by outliers, leading to overfitting on ambiguous features. To solve this issue, we propose a novel loss function:
@@
~$\tilde{\mathcal{L}}(\mathcal{F},\mathcal{G})=\lVert\mathcal{F}-\mathcal{G}\rVert^{\epsilon}\mathcal{L}(\mathcal{F},\mathcal{G}),$ (2)
where $\mathcal{L}$ is from Eq. 1 and $\epsilon$ is a hyper-parameter inspired by the Focal loss [25]. In our experiments, we set $\epsilon=15$.

### v3

@@ Abstract
Three-dimensional local descriptors are crucial for encoding geometric surface properties, making them essential for various point cloud understanding tasks. Among these descriptors, GeDi has demonstrated strong zero-shot 6D pose estimation capabilities but remains computationally impractical for real-world applications due to its expensive inference process. Can we retain GeDi’s effectiveness, while significantly improving its efficiency? In this paper, we explore this question by introducing a knowledge distillation framework that trains an efficient student model to regress local descriptors from a GeDi teacher. Our key contributions include: an efficient large-scale training procedure that ensures robustness to occlusions and partial observations while operating under compute and storage constraints, and a novel loss formulation that handles weak supervision from non-distinctive teacher descriptors. We validate our approach on five BOP Benchmark datasets and demonstrate a significant reduction in inference time while maintaining competitive performance with existing methods, bringing zero-shot 6D pose estimation closer to real-time feasibility. Project website: https://tev-fbk.github.io/dGeDi.
@@ I Introduction
Three-dimensional descriptors (or features) encode geometric surface properties through compact numerical representations [1]. They are essential for various downstream tasks in point cloud understanding, including object 6D pose estimation [3], registration [2], segmentation [4, 5], and navigation [6]. These descriptors can be broadly categorized as either local or global, depending on the type of information they capture. Local descriptors operate at the point or patch level, capturing fine-grained geometric details [7, 8, 2, 9]. Global descriptors also provide point-level representations, but aggregate geometric information from the entire point cloud [10, 11, 12]. Global descriptors are generally more computationally efficient [13] but often sacrifice generalization across different data domains [2].
Local descriptors can achieve rotation invariance through local reference frames (LRF) [2] or point-pair features (PPF) [14]. LRF-based methods establish a local coordinate system for each point using surrounding geometric information [15]. PPF-based methods compute antisymmetric 4D features from pairs of 3D points and their surface normals [16]. Among local descriptors, GeDi [2] has been successfully applied to 6D pose estimation1116D pose estimation aims to recover the 3D position and orientation of an object in a scene from visual sensory data, such as RGBD images. through the FreeZe method [3] in a zero-shot setting, achieving performances that, for the first time in the literature, nearly match those of state-of-the-art fully-supervised methods. However, GeDi’s inference time is impractical for real-world applications due to its computationally expensive process: LRF calculation[15] for each point followed by a PointNet++-based descriptor encoding step [17] (Fig. 1, top).
@@
In this work, we explore how to address this inefficiency by using knowledge distillation [19] to train a student model (Fig. 1, bottom) that regresses GeDi local descriptors, i.e., the teacher model. However, this presents several challenges. Distillation in this context requires large-scale point cloud datasets [20, 21, 22]. The rotation and scale invariance properties of the teacher model have to be effectively transferred to the student model. Low inlier ratios must be handled, as only a small subset of descriptors is typically reliable, resulting in a significant imbalance during distillation, where most points provide unreliable supervision. Despite these challenges, our approach presents unique opportunities. A more efficient descriptor extractor could bring zero-shot 6D pose estimation closer to real-time performance, making it more viable for practical applications. Moreover, by tailoring the distillation process specifically for 6D pose estimation, we can improve descriptor robustness to occlusions and partial observations. To the best of our knowledge, this is the first attempt to apply distillation to 3D local descriptors.
@@
To this end, we introduce a distillation approach to train an efficient neural network (student) based on PointTransformerV3 (PTV3) [18] to generate GeDi descriptors (teacher). We refer to our approach as dGeDi, derived from distilled GeDi. A key challenge in distilling GeDi is balancing efficiency and practicality. On the one hand, a naive online distillation approach, where the teacher is computed jointly with the student, is impractical due to GeDi’s slow run-time, making training excessively slow. On the other hand, a naive offline distillation, where teacher features are precomputed, would be highly storage-inefficient. To overcome these limitations, we propose a novel learning via correspondences paradigm. Offline, we precompute features from a query object, which is the 3D model, while online, we transfer them to the target object, the visible instance of this model present in the scene, using ground-truth transformations from synthetic 6D pose estimation datasets. This allows scalable training on large datasets while keeping compute and storage requirements manageable. Another challenge is that, unlike regression, where supervision comes from precise ground-truth annotations, our distillation supervision can include non-distinctive descriptors (e.g., on the BOP Benchmark [23], only 3% of points can be matched through nearest neighbor search in the teacher feature space). Standard loss functions [24, 25] tend to overfit these outliers, producing unreliable student features. To mitigate this, we introduce a novel loss function that weights each point’s contribution based on its registration error. Points with errors above a certain threshold contribute almost equally, preventing excessive penalization from outliers. Since training data are synthetic, we incorporate data augmentation to enhance rotational robustness in the student model and reduce the synthetic-to-real domain gap. We validate dGeDi through direct comparison with GeDi [2] and further evaluate its effectiveness by integrating it into FreeZe [3], a state-of-the-art zero-shot pose estimator. Additionally, we conduct an extensive ablation study to analyze dGeDi’s key components. Experimental results demonstrate that dGeDi significantly reduces inference time while maintaining competitive performance with alternative methods. In summary, our main contributions are:
@@ II-A Zero-shot object 6D pose estimation
Zero-shot pose estimation methods estimate the 6D pose of objects that are unseen during training. These methods train on synthetic dataset or leverage the knowledge acquired by pre-trained vision and geometric foundational models. The first category of methods like MegaPose [26] and ZeroPose [27] uses render and compare strategy. At first a classification network selects the best-matching template from a set of rendered images that closely resemble the input image. Then, a refiner network iteratively determines the pose by matching the selected template with additional viewpoints. Since these methods requires a separate inference for each input image and template pair, the computational cost and inference time scale linearly with the number of templates. Methods like GigaPose [28] addresses the issue of inference time scaling. It renders templates to extract dense features using a Vision Transformer (ViT), and then finds the best template using fast nearest neighbour search in the feature space. On top, two lightweight MLPs estimate the 2D scale and in plane rotation from a single 2D-2D correspondence and the two relative patch features extracted by another feature extractor. SAM-6D [29] leverages ViT to extract features that are back-projected into 3D to find 3D-3D correspondences between two point clouds. SAM-6D uses a matching network to enhance features before finding the actual correspondences. FoundationPose [30] propose a generalizable and unified pipeline for pose tracking. The coarse pose is estimated using a trained transformer based network. This is followed by a pose refinement network that operates on input image crops and the renderings of the object model. The refined poses are then ranked using a self attention based transformer network.
Other methods do not require task-specific training and leverage the knowledge learned by large scale vision and geometric foundation models. ZS6D [31] extracts features using pre-trained DINOv2 [53] from the inputs image and templates. Then, the best template is chosen using the cosine similarity and 2D-3D correspondences are retrieved using patch-wise similarity. The pose is estimated using PnP and RANSAC. FoundPose [32] integrates DINOv2 patch features into bag-of-words representation for efficient template retrieval. The pose hypothesis is generated using PnP and RANSAC and is refined using optimization-based photometric alignment algorithm. FreeZe [3] fuses descriptors from pre-trained geometric and vision foundation models to extract discriminative features. This is followed by RANSAC-based registration and a symmetry-aware pose refinement.
@@ III Preliminaries: FreeZe method
Preprocessing. Let $\mathcal{P}^{Q}$ and $\mathcal{P}^{T}$ be the 3D point cloud of $Q$ and $T$, respectively. $\mathcal{P}^{Q}$ is obtained by sampling points on $\mathcal{O}$’s surface. $\mathcal{P}^{T}$ is obtained by using an off-the-shelf zero-shot segmentor [29] to extract the crop $\textbf{I}^{T}\subset\textbf{I}$ containing $T$, and then lifting $\textbf{I}^{T}$ in 3D using sensor’s intrinsic parameters.
Feature extraction. FreeZe extracts features $\mathcal{F}^{Q}$, $\mathcal{F}^{T}$ from $\mathcal{P}^{Q},\mathcal{P}^{T}$ using frozen vision [53] and geometric [2] encoders, i.e., $\mathcal{F}=(\mathcal{F}_{\text{geo}}|\mathcal{F}_{\text{vis}})$, where $\mathcal{F}_{\text{geo}}$ are the geometric features, $\mathcal{F}_{\text{vis}}$ are the visual features, and $|$ is the concatenation operator.
@@ IV Object-oriented distillation
While FreeZe performs feature extraction with both geometric and vision encoders, we focus on the distillation of the former one [^2] Object-oriented distillation can also be applied to the vision encoder, but it is out of scope for this work., assuming $\mathcal{F}=\mathcal{F}_{\text{geo}}$. We introduce a novel distillation procedure to transfer the distinctive properties of a slow teacher encoder to a faster student encoder (Fig. 2). We define our proposed distillation approach as object-oriented because training supervision is provided at the object level. First, we extract teacher descriptors $\mathcal{F}^{Q}=\Phi_{\Theta}(\mathcal{P}^{Q}),\mathcal{F}^{T}=\Phi_{\Theta}(\mathcal{P}^{T})$ using GeDi [2] encoder $\Phi_{\Theta}$. Then, we learn student descriptors $\mathcal{G}^{Q}=\Psi_{\Omega}(\mathcal{P}^{Q}),\mathcal{G}^{T}=\Psi_{\Omega}(\mathcal{P}^{T})$ with a PTV3 [18] encoder $\Psi_{\Omega}$. During distillation, we optimize the parameters $\Omega$ so that $\mathcal{G}^{Q}\approx\mathcal{F}^{Q}$ and $\mathcal{G}^{T}\approx\mathcal{F}^{T}$, while $\Theta$ remain frozen. Note that this differs from online knowledge distillation approaches [56], where both the teacher and student are neural networks and learn $\Theta,\Omega$ simultaneously. In our case, the architectures of $\Phi_{\Theta}$ and $\Psi_{\Omega}$ are distinct (Fig. 1). Teacher features are precomputed, stored in memory, and loaded as needed during distillation. To optimize this process, we propose storing teacher features only for query objects and introduce a module that leverages ground-truth 6D poses to transfer them to target objects (Sec. IV-A). Moreover, we propose a custom loss function that focuses learning on noise-free points, leading to improved performance (Sec. IV-B).
@@ IV-A Learning via correspondences
During training, given a data sample $(\mathcal{P}^{Q},\mathcal{P}^{T},\textbf{T}^{Q\to T})$, we leverage the ground-truth 6D pose $\textbf{T}^{Q\to T}$ to align $\mathcal{P}^{Q}$ with $\mathcal{P}^{T}$ and perform a nearest neighbor search in the metric space to establish the correspondence $\Gamma\colon\mathcal{P}^{T}\to\mathcal{P}^{Q}$, where each point $\textbf{p}\in\mathcal{P}^{T}$ is mapped to $\textbf{q}=\operatorname*{argmin}_{\textbf{q}\in\mathcal{P}^{Q}}\lVert\textbf{T}^{Q\to T}\textbf{q}-\textbf{p}\rVert$, where p and q represent the 3D coordinates of corresponding points. We use $\Gamma$ to transfer query teacher descriptors $\mathcal{F}^{Q}=\Phi_{\Theta}(\mathcal{P}^{Q})$ to $\mathcal{P}^{T}$, obtaining $\tilde{\mathcal{F}}^{T}=\Gamma(\mathcal{F}^{Q})$. By replacing the explicit computation of $\mathcal{F}^{T}=\Phi_{\Theta}(\mathcal{P}^{T})$ with $\tilde{\mathcal{F}}^{T}$ during distillation, we achieve two key benefits: (i) improved supervision for target objects and (ii) a scalable, object-oriented distillation process.
Improved target supervision. Matching features $\mathcal{F}^{Q},\mathcal{F}^{T}$ is challenging because input data differ significantly: $\mathcal{P}^{Q}$ is computed from noise-free synthetic CAD models, while $\mathcal{P}^{T}$ is extracted from noisy real-world RGBD images. This domain gap results in variations in point densities, partiality, and artifacts created by lifting the input image I to generate $\mathcal{P}^{T}$. For example, on the MegaPose-GSO [26] dataset, we measured that, on average, only 21.4% of the points can be matched via nearest neighbor search in the teacher feature space. This means that 78.6% of the features are considered outliers for feature matching. In contrast, by transferring query teacher features to target point clouds with $\Gamma$, we simulate the ideal scenario where all target points are matched with the corresponding query points, promoting reliable supervision.
@@ IV-B Loss function design
Although GeDi’s descriptors have proven to be both generalizable and distinctive [2], they still exhibits low inlier ratios, with an average RON [6] of 3% on the BOP Benchmark and 21.4% on MegaPose-GSO. This indicates that only a small fraction of points are highly distinctive, while the majority have ambiguous features. Consequently, a loss function that assigns equal weight to all points will be dominated by outliers, leading to overfitting on ambiguous features. To solve this issue, we propose a novel loss function:
@@
$\tilde{\mathcal{L}}(\mathcal{F},\mathcal{G})=\lVert\mathcal{F}-\mathcal{G}\rVert^{\epsilon}\mathcal{L}(\mathcal{F},\mathcal{G}),$ (2)
where $\mathcal{L}$ is from Eq. 1 and $\epsilon$ is a hyper-parameter inspired by the Focal loss [25]. In our experiments, we set $\epsilon=15$ to guarantee that the points with large regression error receive significantly higher weights than the points with small error.